Grade 12th passAlgebra

Simplify and express each of the following in exponential form:
(23 × 2)2

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

= (23 + 1)2 ... [∵am × an = am + n]
= (24)2
(24)2 can be written as = (2)4 x 2 ... [∵(am)n = amn]
= 28

​Thanks